what is the slope intercept equation for the line with the given slope that contains the points
m=-3;(1,-5)
I know that the answer is -3-2 but how do I reach that answer?

Answer by dabanfield(803)

what is the slope intercept equation for the line with the given slope that contains the points
m=-3;(1,-5)
I know that the answer is -3-2 but how do I reach that answer?
The slope intercept form of a line is y = m*x + b where m is the slope and b is the value of y where the line intercepts the y-axis (that is when x=0).
Since m = -3 we have:
1.) y = -3*x + b.
The line passes through the point (1,-5) so when x = 1, y = -5. Substituting these values for x and y in equation 1.) we have:
-5 = -3*1 + b
-5 = -3 + b
b = -5 + 3 = -2
With b = -2 equation 1.) becomes:
y = -3x - 2